Given: G = 6.67259 × 10−11 Nm2/kg2 .
A 1340 kg geosynchronous satellite or-
bits a planet similar to Earth at a radius
1.97 × 105 km from the planet’s center. Its
angular speed at this radius is the same as the
rotational speed of the Earth, and so they ap-
pear stationary in the sky. That is, the period
of the satellite is 24 h .
What is the force acting on this satellite?
Answer in units of N
012 (part 2 of 2) 10.0 points
What is the mass of this planet?
Answer in units of kg
1
Expert's answer
2013-02-19T07:54:22-0500
1) The forceacting on this satellite:
F = G*m*M/R^2
G - gravitational constant
m - mass of thу satellite
M - mass of the planet
R - radius of the orbit
M = (2Pi/T)^2/G*R^3 (from 2nd part)
F = G*m*(2Pi/T)^2/G*R^3/R^2 = m*(2Pi/T)^2*R = 1396.06 H
2) Newton's second laws of motion:
mV^2/R = G*m*M/R^2
V - speed of the satellite.
V^2 = G*M/R
M = R*V^2/G
2*Pi*R/T = V
M = R*(2*Pi*R/T)^2/G = (2Pi/T)^2/G*R^3 = 6.0595*10^26 kg